# A Scalable Matrix Visualization for Understanding Tree Ensemble Classifiers

### v1

@@ I Introduction
Second, simply using model reduction inevitably results in a loss of fidelity to the original model. This causes discrepancies between the behavior of the reduced model and that of the original one. Therefore, only examining the reduced rule set after model reduction may lead to problems in analyzing the whole model. For example, an abnormal behavior identified in the reduced model may be difficult to trace back to rules in the original model without access to the whole rule set. To address this problem, it is essential to find an effective way to examine all the rules of the original model. Hierarchical visualization has been shown effective in the exploration of large-scale data [36]. By re-organizing rather than discarding the rules that are not preserved in model reduction, the issue of losing fidelity can be effectively mitigated. However, the pre-built hierarchies commonly used in hierarchical visualization methods [8, 56, 60] lack the flexibility to accommodate different user preferences during exploratory analysis. This flexibility is often needed in real-world applications. For example, in the credit card approval process, users frequently examine different subsets of rules by constraining the values of different attributes (e.g., Income, Years Employed, etc.) in different ranges. As a result, there is a need to flexibly build the hierarchy for different analysis needs.
@@ II Related Work
Figure 1: An illustration of a tree ensemble classifier.
@@ II-A Model Surrogate
Our model reduction method relates to a category of explanation methods that utilize surrogate models. These methods use a simpler and more interpretable model to approximate and explain complex models [12, 43, 50, 34, 59].
Most model surrogate methods are model-agnostic. These methods derive a surrogate model from any given machine learning model by approximating the input-output relationship of the model. While these methods have good generalizability, they inevitably fail to take into account the inner decision logic such as node splitting and majority voting in ensemble tree models. As a result, these methods cannot precisely reveal how predictions are made.
For tree ensemble models, the surrogate methods proposed in [32, 53] leverage the characteristics of tree ensembles for more accurate approximations. Meinshausen [32] selected representative rules from a random forest by solving a quadratic programming problem with linear inequality constraints. Vidal et al. [53] constructed an equivalent single decision tree from a given tree ensemble to improve model interpretability. Although both methods successfully preserve common rules that summarize the behavior of the original model well, they ignore anomalous rules that are important for revealing potential flaws in the model. In contrast, our method re-organizes a tree ensemble into a hierarchy of rules to mitigate the issues of losing fidelity, and uses anomaly-biased model reduction to maintain a balance between the common rules and the anomalous rules at each hierarchical level.
@@ II-B Visualization for Model Analysis
The visualization community has developed a number of methods and systems to visually interpret, analyze, and diagnose machine learning models. Initial efforts and their subsequent variations aim to illustrate a variety of performance measures such as precision, recall, and between-class confusion [42, 21]. Most recent efforts focus on explaining the inner workings of a machine learning model based on architecture exploration and the reasoning of its prediction.
Architecture visualization methods focus on explaining the inner workings of the model by analyzing specific components in the model, such as neurons and kernels in convolutional neural networks [51, 30, 4], hidden states in recurrent neural networks [33, 48, 6, 28, 45], attention heads in transformers and other attention-based models [47, 14], or tree structures in decision trees [52, 35, 31]. These visualizations are designed to help model developers analyze and diagnose machine learning models. However, for domain experts who have little knowledge of machine learning, such as financial analysts or doctors, the complex workflows and model-architecture-specific designs are of little help for their understanding and utilizing the models. To support these domain experts, our work focuses on explaining rules learned from tree ensemble classifiers, as rules are more understandable to them. If offered the right tool, they can make sense of the learned rules and validate the rules of interest.
@@ II-C Rule Visualization
The node-link diagram [52, 49, 35, 26] naturally represents the decision tree structure or summarizes multiple rules by aggregating the same conditional statement. For example, BaobabView [52] employs a node-link diagram to visualize the rules of decision trees and further encodes the split points and data flows in the tree. TreePOD [35] extends this design by utilizing a treemap to display the distribution of samples covered by different tree nodes. SuRE [59] summarizes the entire random forest into a node-link diagram while also retaining useful logical relationships between different rules. While the node-link diagrams effectively illustrate relationships between rules, they struggle to remain clear and manageable when displaying tens of thousands of rules. The matrix diagram [34, 37, 61] is a natural graphical representation of a list of rules. RuleMatrix [34] visualizes a falling rule list using a matrix diagram with rows as rules, columns as attributes, and cells as conditions using histogram encoding. ExplainableMatrix [37] also employs a matrix diagram, but distinguishes itself by using colored rectangles in matrix cells to encode the conditions for different classes. However, these methods show all the rules in one view without filtering, and still encounter the scalability issue in actual scenarios.
@@ III-A Design Goals
R2.2 Cater to different analysis preferences. Users may have different analysis needs and focus on different subsets of rules during the exploration. For example, E7 mentioned that in the overall analysis of stocks, he would pay more attention to the attributes directly related to price changes, such as moving average price and price volatility. When he focused on low-cap stocks, he would like to examine the attributes related to stock liquidity, such as turnover ratio and market depth. With the different analysis needs, he zoomed into different subsets of rules. Compared to a static hierarchy, which organizes the rules with a pre-defined criterion, a dynamic hierarchy is better at catering to users’ different preferences [18, 9].
@@ IV-B Model Reduction
For a specific sample $x_{i}$, as shown in Fig. 3, the multi-class hinge loss encourages the score for class $\ell_{i}$ to be greater than any other class by a margin of at least $\xi$. This helps improve the generalization ability of the extracted rules [44]. In our implementation, the parameters $\xi$ and $\lambda$ are simultaneously determined using grid search to strike a balance between fidelity and anomaly detection. $M$ is set to 80 based on the screen constraints and suggestions from the experts.
Eq. (1) is an integer linear programming problem, which is NP-hard and time-consuming to solve. To solve it efficiently, we relax the $z_{j}$ to continuous variables. The relaxed problem becomes a linear programming problem and can be solved in polynomial time [25]. Then, we adopt deterministic rounding to discretize $z_{j}$ and determine which rules belong to $R^{*}$.
@@ V-C2 Analyzing rules of interest
Here, $\pi(i)$ is the index of the rule in $R$ which is reordered to row $i$ in the matrix. The first term is an indicator function, which equals $1$ when the predictions of the two rules are consistent, and 00 otherwise. The second term $\text{sim}_{j}(r_{\pi(i)},r_{\pi(i+1)})$ is also an indicator function that determines whether the two rules have similar conditions on the $j$th attribute. For categorical attributes, it is defined as whether the conditions are the same. For numerical attributes, it is defined as whether both the left and right endpoints between the two conditions differ within a user-specified threshold $\tau$. We set $\tau=0.1$ by default, which gives satisfactory results in practice.
Maximizing $S_{j}$ is equivalent to solving a traveling salesman problem, which will reorder the rules and form rule groups on the attribute $A_{j}$. Fig. 7 shows examples of rule groups formed on each attribute. However, reordering rules to optimize all $S_{j}$ simultaneously is impossible as different optimization goals may conflict with each other. As experts are more concerned with patterns on important/preferred attributes, we adopt the idea of the constraint method [13], which maximizes $S_{1},S_{2},\ldots,S_{m}$ in turn. The constraints are that when maximizing $S_{j}$, the reordering of rules in a formed rule group can only be carried out within the group without crossing the group boundaries. This ensures minimal changes to $S_{1},S_{2},\ldots,S_{j-1}$ when maximizing $S_{j}$. We use the divide-and-conquer strategy proposed by Liu et al. [30] to solve this constrained traveling salesman problem.
@@ VI Evaluation
Dataset #Classes #Samples Model Ours RuleMatrix HSR Node harvest
@@
Credit Card 2 690 RF 0.9538 0.1309 0.9503 0.0880 0.9264 0.0000 0.9300 0.0473
GBT 0.9249 0.1271 0.9075 0.0735 0.8682 0.0846 0.9017 0.0042
Wine Quality 2 1,599 RF 0.8725 0.1432 0.8665 0.0689 0.7484 0.0000 0.5175 0.1660
GBT 0.8350 0.0974 0.8065 0.0329 0.7368 0.0000 0.8275 0.0299
Crime 2 1,994 RF 0.9178 0.1999 0.9042 0.0203 0.8910 0.0000 0.8577 0.1602
GBT 0.9118 0.1152 0.8986 0.0119 0.8368 0.0000 0.8978 0.0489
Abalone 4 4,177 RF 0.8392 0.1247 0.8243 0.0085 0.6695 0.0081 N/A N/A
GBT 0.8967 0.1263 0.8268 0.1231 0.5715 0.3553 N/A N/A
Obesity 7 2,111 RF 0.8561 0.1732 0.7720 0.0138 0.7277 0.0683 N/A N/A
GBT 0.9470 0.2309 0.8394 0.0004 0.7233 0.0084 N/A N/A
Dry Bean 7 13,611 RF 0.9738 0.2829 0.9702 0.0032 0.8726 0.0193 N/A N/A
GBT 0.9151 0.3032 0.9151 0.0946 0.8369 0.2360 N/A N/A
@@ VI-A Quantitative Experiment
Datasets. We conducted the quantitative experiment on six datasets from the UCI Machine Learning Repository [2], as used in previous works [34, 59], including three binary classification datasets (Credit Card, Wine Quality, Crime) and three multiple classification ones (Abalone, Obesity, Dry Bean).
@@
Experiment settings. To demonstrate the generalizability of our method to different tree ensemble classifiers, we conducted experiments on random forest models and gradient boosted tree models. Each dataset was partitioned into a 75% training set and a 25% test set. We first trained each model on the training set, resulting in models with rule counts ranging from 2,126 (LightGBM on credit card dataset) to 52,455 (LightGBM on drybean dataset). Then, we performed model reduction to approximate the predictions on the training set. All methods are required to select 80 rules (consistent with Sec. IV) for this approximation. The fidelity was calculated on the test set, and the average anomaly score was calculated on the generated rule set. We ran five trials for each setting to reduce the effect of randomness.
@@ VI-B1 Credit Card Approval
In this case study, we collaborated with E1 and E2 to illustrate how RuleExplorer improves the understanding and diagnosis of the model to approve credit card applications. E1 is a bank account manager with two years of work experience, who aims to increase credit card approval rates while minimizing the risk of defaults. E2 is a data scientist responsible for developing the credit card approval model and improving its performance. Both E1 and E2 have participated in the interview for collecting design goals in Sec. III. In this case study, they used the Australian Credit Approval dataset [41]. The dataset consists of 690 samples, where each sample represents a credit card applicant, with 16 attributes and a label indicating approval or rejection. The dataset includes 307 “approved” samples and 383 “rejected” samples. These samples were randomly divided into training and test data with a ratio of 75% and 25%, respectively. To comprehensively evaluate the model, E2 used Synthetic Data Vault [39] to augment the test data based on its distribution and labeled the new test samples in collaboration with E1. After the augmentation, there are 500 samples in the test data. A random forest model with 200 trees and a maximum depth of 10 has been trained on this dataset with the scikit-learn library [40]. The trained model includes 7,085 rules and achieves an accuracy of 83.60%.
@@
Analyzing common rules. By default, the rules are sorted in descending order of their coverage. E1 first analyzed common rules (rules with higher coverage) and found that most of them used the attribute PriorDefault (Fig. 5B). He wanted to examine how PriorDefault influenced the model predictions. To this end, he clicked on the header of PriorDefault, and the rules were grouped based on their conditions on this attribute. There are three groups. (Fig. 8). At the top, the rules with the condition of no prior defaults made predictions of “rejected” (Fig. 8A). The matrix reordering algorithm reveals several patterns, such as low income, few years employed, or low credit scores, which accounted for the rejection. In the middle, the rules with the condition of no prior defaults made predictions of “approved” (Fig. 8B). Compared to the group at the top, these rules had additional positive conditions, such as a certain income level or being older, indicating a lower risk of default and a more stable status. E1 regarded the behavior (Fig. 8A and B) as reasonable, since the condition of no prior defaults cannot directly lead to the approval, and it should be combined with additional conditions for further consideration. At the bottom, the rules with the condition of prior defaults consistently led to predictions of “rejected” (Fig. 8C). E1 selected these rules and zoomed into their neighborhood in the dynamic hierarchy. As a result, he found that all these rules resulted in “rejected” (Fig. 8D). Since there are some real-world cases where applicants with prior default records are still approved, E1 was curious why this was not reflected in these rules, and consulted with E2. E2 explained that a credit card approval model typically rejects all applicants with prior default records in realistic scenarios and leaves disputed cases for manual processing. E1 agreed with this strategy, as applicants with default records pose a higher credit risk. To summarize, E1 was satisfied with the behavior of the rules utilizing PriorDefault. Similarly, E1 analyzed other frequently used attributes, including Job, Income, Years Employed, and acknowledged that these attributes were properly employed to make predictions.
@@
E1 continued to examine common rules and noticed a rule that had significantly lower confidence than the rules with similar coverage (Fig. 5C). The low confidence indicated that this rule covered many mispredicted samples. Therefore, he examined its covered samples. He found that these samples shared some negative conditions (Fig. 5D), including unemployment, missing credit scores, lack of income, and having prior default records. Although these samples were labeled as “approved” (Fig. 5E), E1 suspected they should be rejected given those negative conditions. He then filtered the samples with such negative conditions and found that they were mostly labeled as “rejected.” This indicated that there were inconsistent approval criteria in the training data. He raised this issue with E2, and E2 commented that such inconsistency harmed the model training and led to performance degradation. Therefore, E1 unified the approval criteria in these samples and corrected their labels to be “rejected.” By retraining the model on the corrected data, the accuracy increased from 83.60% to 86.20%, and the prediction confidence increased from 0.7403 to 0.7423.
@@
Analyzing anomalous rules. To identify potential flaws in the model, E1 turned to analyze the anomalous rules. He sorted the rules according to their anomaly scores. The low coverage of the anomalous rules suggested that they were learned from only a few samples. E1 started his analysis by examining the rule with the highest anomaly score. As shown in Fig. 9A1, this rule suggests approval for young, single, unemployed applicants with a medium-to-low income. These conditions were mostly neutral or negative, which should not lead to an approval. He zoomed in to examine its neighborhood and observed that they shared few conditions (Fig. 9B1), which also indicated that its behavior differed a lot from other rules. To analyze why this rule was learned, he clicked on it and examined its covered samples. As shown in Fig. 9C1, it covered only an 18-year-old applicant with a 10-year work experience (#321). It is suspicious that an 18-year-old has a 10-year work experience. E1 confirmed that this applicant should not be approved and corrected its label to “rejected”.
Subsequently, E1 examined the rule with the second-highest anomaly score. This rule suggested approval for applicants with some debt, having a low credit score, having a driver’s license, and having a specific job of type #7 (Fig. 9A2). As these conditions were primarily neutral or negative, E1 doubted the approval decision for these applicants. Therefore, he checked the samples covered by this rule and realized that most of the applicants had no prior default records and a certain level of income (Fig. 9B2). E1 believed these positive conditions were the real reason for the approval. By filtering samples with such positive conditions using the scented widgets in the attribute view (Fig. 9C2), he found that most of them (95.55%) are approved (Fig. 9D2). Thus, he suspected that this rule occurred because of the spurious association between the irrelevant conditions and the approval decision in these covered samples (Fig. 9A2). E1 then discussed this issue with E2. E2 considered adding more diverse samples, especially those that meet the conditions of this rule but get rejected, would mitigate this spurious association. He thus generated more samples using data augmentation, and E1 validated these samples to ensure they accurately reflected real-world scenarios.

### v2

# RuleExplorer: A Scalable Matrix Visualization for Understanding Tree Ensemble Classifiers

@@ I Introduction
Second, simply using model reduction inevitably results in a loss of fidelity to the original model. This causes discrepancies between the behavior of the reduced model and that of the original one. Therefore, only examining the reduced rule set after model reduction may lead to problems in analyzing the whole model. For example, an abnormal behavior identified in the reduced model may be difficult to trace back to rules in the original model without access to the whole rule set. To address this problem, it is essential to find an effective way to examine all the rules of the original model. Hierarchical visualization has been shown effective in the exploration of large-scale data [42]. By re-organizing rather than discarding the rules that are not preserved in model reduction, the issue of losing fidelity can be effectively mitigated. However, the pre-built hierarchies commonly used in hierarchical visualization methods [37, 11, 62, 67, 13] lack the flexibility to accommodate different user preferences during exploratory analysis. This flexibility is often needed in real-world applications. For example, in the credit card approval process, users frequently examine different subsets of rules by constraining the values of different attributes (e.g., Income, Years Employed, etc.) in different ranges. As a result, there is a need to flexibly build the hierarchy for different analysis needs.
@@ II Related Work
Figure 1: An illustration of a tree ensemble classifier used to predict stock price movement with labels “increase”, “decrease”, and “stable.”
@@ II-A Model Surrogate
Our model reduction method relates to a category of explanation methods that utilize surrogate models. These methods use a simpler and more interpretable model to approximate and explain complex models [17, 49, 56, 40, 66, 10].
Most model surrogate methods are model-agnostic. These methods derive a surrogate model from any given machine learning model by approximating the input-output relationship of the model. While these methods have good generalizability, they inevitably neglect the inner decision logic in tree ensemble models, such as node splitting and majority voting. As a result, these methods cannot precisely reveal how predictions are made.
For tree ensemble models, the surrogate methods proposed in [38, 59] leverage the characteristics of tree ensembles for more accurate approximations. Meinshausen [38] selected representative rules from a random forest by solving a quadratic programming problem with linear inequality constraints. Vidal et al. [59] constructed an equivalent single decision tree from a given tree ensemble to improve model interpretability. Although these methods successfully preserve common rules that well summarize the behavior of the original model, they inevitably lose fidelity because other rules are discarded, especially those anomalous rules that are important for revealing potential flaws in the model. In contrast, our method re-organizes a tree ensemble into a hierarchy of rules to mitigate the issues of losing fidelity, and uses anomaly-biased model reduction to maintain a balance between the common rules and the anomalous rules at each hierarchical level.
@@ II-B Visualization for Model Analysis
The visualization community has developed a number of methods and systems to visually interpret, analyze, and diagnose machine learning models. Initial efforts and their subsequent variations aim to illustrate a variety of performance measures such as precision, recall, and between-class confusion [48, 26, 12]. Most recent efforts focus on explaining the inner workings of a machine learning model based on architecture exploration and the reasoning of its prediction.
Architecture visualization methods focus on explaining the inner workings of the model by analyzing specific components in the model [63], such as neurons and kernels in convolutional neural networks [57, 35, 5], hidden states in recurrent neural networks [39, 54, 8, 33, 51], attention heads in transformers and other attention-based models [53, 19], or tree structures in decision trees [58, 41, 36]. These visualizations are designed to help model developers analyze and diagnose machine learning models. However, for domain experts who have little knowledge of machine learning, such as financial analysts or doctors, the complex workflows and model-architecture-specific designs are of little help for their understanding and utilizing the models. To support these domain experts, our work focuses on explaining rules learned from tree ensemble classifiers, as rules are more understandable to them. If offered the right tool, they can make sense of the learned rules and validate the rules of interest.
@@ II-C Rule Visualization
The node-link diagram [58, 55, 41, 31] naturally represents the decision tree structure or summarizes multiple rules by aggregating the same conditional statement. For example, BaobabView [58] employs a node-link diagram to visualize the rules of decision trees and further encodes the split points and data flows in the tree. TreePOD [41] extends this design by utilizing a treemap to display the distribution of samples covered by different tree nodes. SuRE [66] summarizes the entire random forest into a node-link diagram while also retaining useful logical relationships between different rules. While the node-link diagrams effectively illustrate relationships between rules, they struggle to remain clear and manageable when displaying tens of thousands of rules. The matrix diagram [40, 43, 68] is a natural graphical representation of a list of rules. RuleMatrix [40] visualizes a falling rule list using a matrix diagram with rows as rules, columns as attributes, and cells as conditions using histogram encoding. DeforestVis [10] provides attribute-based explanations using simplified decision trees, allowing users to incrementally generate them to explore the trade-off between complexity and fidelity. ExplainableMatrix [43] also employs a matrix diagram, but distinguishes itself by using colored rectangles in matrix cells to encode the conditions for different classes. However, these methods show all the rules in one view without filtering, and still encounter the scalability issue in actual scenarios.
@@ III-A Design Goals
R2.2 Cater to different analysis preferences. Users may have different analysis needs and focus on different subsets of rules during the exploration. For example, E7 mentioned that in the overall analysis of stocks, he would pay more attention to the attributes directly related to price changes, such as moving average price and price volatility. When he focused on low-cap stocks, he would like to examine the attributes related to stock liquidity, such as turnover ratio and market depth. With the different analysis needs, he zoomed into different subsets of rules. Compared to a static hierarchy, which organizes the rules with a pre-defined criterion, a dynamic hierarchy is better at catering to users’ different preferences [23, 14, 71].
@@ IV-B Model Reduction
For a specific sample $x_{i}$, as shown in Fig. 3, the multi-class hinge loss encourages the score for class $\ell_{i}$ to be greater than any other class by a margin of at least $\xi$. This helps improve the generalization ability of the extracted rules [50]. To reduce the user burden of parameter adjustments and improve ease of use, we determine $\xi$ and $\lambda$ using a grid search. This search spans values from from 0.1 to 1.0, with a step size of 0.1, effectively balancing fidelity and anomaly scores. Specifically, we first set $\lambda$ to 0 and searched for $\xi$ that achieved the highest fidelity, then adjusted $\lambda$ to improve anomaly scores while keeping fidelity loss within 1%. $M$ is set to 80 based on the screen constraints and suggestions from the experts, as they prefer to examine as many rules as possible while maintaining visual clarity for rule analysis. We also conducted a sensitivity analysis of $M$ in the supplemental materials to justify this choice.
Eq. (1) is an integer linear programming problem, which is NP-hard and time-consuming to solve. Even advanced commercial solvers cannot solve it within reasonable time limits due to the extremely large basis matrix and overwhelming memory consumption resulting from thousands of rules and samples. To solve it efficiently, we relax the $z_{j}$ to continuous variables. The relaxed problem becomes a linear programming problem and can be solved in polynomial time [30]. Then, we adopt deterministic rounding to discretize $z_{j}$ and determine which rules belong to $R^{*}$.
@@ V-C2 Analyzing rules of interest
Here, $\pi(i)$ is the index of the rule in $R$ which is reordered to row $i$ in the matrix. The first term is an indicator function, which equals $1$ when the predictions of the two rules are consistent, and 00 otherwise. The second term $\text{sim}_{j}(r_{\pi(i)},r_{\pi(i+1)})$ is also an indicator function that determines whether the two rules have similar conditions on the $j$th attribute. For categorical attributes, it is defined as whether the conditions are the same. For numerical attributes, it is defined as whether both the left and right endpoints between the two conditions differ within a user-specified threshold $\tau$. We set $\tau=0.1$ by default. This threshold controls whether two numerical conditions are similar enough to be placed adjacently. A smaller $\tau$ results in stricter matching criteria, while a larger $\tau$ value relaxes the criteria.
Maximizing $S_{j}$ is equivalent to solving a traveling salesman problem [2], which will reorder the rules and form rule groups on the attribute $A_{j}$. Fig. 7 shows examples of rule groups formed on each attribute. However, reordering rules to optimize all $S_{j}$ simultaneously is impossible as different optimization goals may conflict with each other. As experts are more concerned with patterns on important/preferred attributes, we adopt the idea of the constraint method [18], which maximizes $S_{1},S_{2},\ldots,S_{m}$ in turn. The constraints are that when maximizing $S_{j}$, the reordering of rules in a formed rule group can only be carried out within the group without crossing the group boundaries. This ensures minimal changes to $S_{1},S_{2},\ldots,S_{j-1}$ when maximizing $S_{j}$. We use the divide-and-conquer strategy proposed by Liu et al. [35] to solve this constrained traveling salesman problem.
@@ VI Evaluation
Dataset #Classes #Attributes #Samples Model #Rules Ours RuleMatrix HSR Node harvest
@@
Credit Card 2 14 690 RF 7,085 0.9538 0.1309 0.9503 0.0880 0.9264 0.0000 0.9300 0.0473
GBT 2,126 0.9249 0.1271 0.9075 0.0735 0.8682 0.0846 0.9017 0.0042
Wine Quality 2 11 1,599 RF 28,116 0.8725 0.1432 0.8665 0.0689 0.7484 0.0000 0.5175 0.1660
GBT 7,474 0.8350 0.0974 0.8065 0.0329 0.7368 0.0000 0.8275 0.0299
Crime 2 127 1,994 RF 12,576 0.9178 0.1999 0.9042 0.0203 0.8910 0.0000 0.8577 0.1602
GBT 4,322 0.9118 0.1152 0.8986 0.0119 0.8368 0.0000 0.8978 0.0489
Abalone 4 8 4,177 RF 23,500 0.8392 0.1247 0.8243 0.0085 0.6695 0.0081 N/A N/A
GBT 11,810 0.8967 0.1263 0.8268 0.1231 0.5715 0.3553 N/A N/A
Obesity 7 16 2,111 RF 19,490 0.8561 0.1732 0.7720 0.0138 0.7277 0.0683 N/A N/A
GBT 30,458 0.9470 0.2309 0.8394 0.0004 0.7233 0.0084 N/A N/A
Dry Bean 7 16 13,611 RF 6,209 0.9738 0.2829 0.9702 0.0032 0.8726 0.0193 N/A N/A
GBT 52,455 0.9151 0.3032 0.9151 0.0946 0.8369 0.2360 N/A N/A
@@ VI-A Quantitative Experiment
Datasets. We conducted the quantitative experiment on six datasets from the UCI Machine Learning Repository [3], as used in previous works [40, 66], including three binary classification datasets (Credit Card, Wine Quality, Crime) and three multiple classification ones (Abalone, Obesity, Dry Bean). The attribute numbers across these datasets range from 8 to 127, while the sample numbers range from 690 to 13,611. The diverse characteristics of these datasets enable a robust evaluation across various tasks and dataset scales.
@@
Experiment settings. To demonstrate the generalizability of our method to different tree ensemble classifiers, we conducted experiments on random forest models and gradient boosted tree models. Each dataset was partitioned into a 75% training set and a 25% test set. We first trained each model on the training set, resulting in models with rule counts ranging from 2,126 (LightGBM on credit card dataset) to 52,455 (LightGBM on drybean dataset). This wide range of rule numbers ensures a thorough evaluation across different levels of model complexity. Then, we performed model reduction to approximate the predictions on the training set. All methods are required to select 80 rules (consistent with Sec. IV) for this approximation. The fidelity was calculated on the test set, and the average anomaly score was calculated on the generated rule set. We ran five trials for each setting to reduce the effect of randomness.
@@ VI-B1 Credit Card Approval
In this case study, we collaborated with E1 and E2 to illustrate how RuleExplorer improves the understanding and diagnosis of the model to approve credit card applications. E1 is a bank account manager with two years of work experience, who aims to increase credit card approval rates while minimizing the risk of defaults. E2 is a data scientist responsible for developing the credit card approval model and improving its performance. Both E1 and E2 have participated in the interview for collecting design goals in Sec. III. During the study, E1 interacted with the system to explore and identify potential model issues, while E2 helped explain the underlying causes of these findings from a machine learning perspective and suggested ways to improve the model. The study used the Australian Credit Approval dataset [47], which consists of 690 samples, each representing a credit card applicant with 16 attributes and a label indicating approval or rejection. The dataset includes 307 “approved” samples and 383 “rejected” samples. These samples were randomly divided into training and test data with a ratio of 75% and 25%, respectively. To comprehensively evaluate the model, E2 used Synthetic Data Vault [45] to augment the test data based on its distribution and labeled the new test samples in collaboration with E1. The detailed augmentation process can be found in the supplemental material. After the augmentation, there are 500 samples in the test data. A random forest model with 200 trees and a maximum depth of 10 has been trained on this dataset with the scikit-learn library [46]. The trained model includes 7,085 rules and achieves an accuracy of 83.60%.
@@
Analyzing common rules. By default, the rules are sorted in descending order of their coverage, which facilitates the analysis of common rules (rules with higher coverage). E1 found that most of the common rules used the attribute PriorDefault (Fig. 5B). He wanted to examine how this attribute influenced the model predictions. To this end, he clicked on the header of PriorDefault, and the matrix reordering algorithm grouped the rules based on their predictions and their conditions on this and subsequent attributes. There are three groups (Fig. 8). Rules with the condition of no prior defaults but predictions of “rejected” (Fig. 8A) linked to negative conditions such as low income, few years employed, or low credit scores. Rules with the condition of no prior defaults and predictions of “approved” (Fig. 8B) had additional positive conditions, such as a certain income level or being older, indicating a lower risk of default and a more stable status. E1 regarded the behavior (Fig. 8A and B) as reasonable, since the condition of no prior defaults cannot directly lead to the approval, and it should be combined with additional conditions for further consideration. The third group included rules with the condition of prior defaults and predictions of “rejected” (Fig. 8C). Since there are some real-world cases where applicants with prior default records are still approved, E1 wanted to see whether these were reflected. He selected the rules in (Fig. 8C) and zoomed into their neighborhood in the dynamic hierarchy, and found that all rules in the neighborhood had predictions of “rejected” (Fig. 8D). Consultation with E2 clarified that a credit card approval model typically rejects all applicants with prior default records in realistic scenarios and leaves disputed cases for manual processing. To summarize, E1 was satisfied with the behavior of the rules utilizing PriorDefault. Similarly, E1 analyzed other frequently used attributes, including Job, Income, Years Employed, and acknowledged their proper behavior in making predictions.
@@
E1 continued to examine common rules and noticed a rule that had significantly lower confidence than the rules with similar coverage (Fig. 5C). The low confidence indicated that this rule covered many mispredicted samples. The data table enabled him to examine the samples covered by this rule. He found that these samples shared some negative conditions (Fig. 5D), including unemployment, missing credit scores, lack of income, and having prior default records, but were labeled as “approved” (Fig. 5E). E1 suspected that they should be rejected given those negative conditions. He then used the attribute view to filter the samples with such negative conditions and found in the info view that they were mostly labeled as “rejected.” This indicated that there were inconsistent approval criteria in the training data. He raised this issue with E2, and E2 commented that such inconsistency harmed the model training and led to performance degradation. Therefore, E1 unified the approval criteria in these samples and corrected their labels to be “rejected.” By retraining the model on the corrected data, the accuracy increased from 83.60% to 86.20%, and the prediction confidence increased from 0.7403 to 0.7423.
@@
Analyzing anomalous rules. Reordering rules according to their anomaly scores facilitates the analysis of anomalous rules. E1 started his analysis by examining the rule with the highest anomaly score. As shown in Fig. 9A1, this rule suggests approval for young, single, unemployed applicants with a medium-to-low income. These conditions were mostly neutral or negative, which should not lead to approval. He zoomed into its neighborhood and observed that few conditions were shared (Fig. 9B1), which also indicated that its behavior differed a lot from other rules. He then examined its covered samples in the data table. As shown in Fig. 9C1, it covered only an 18-year-old applicant with a 10-year work experience (#321), which was suspicious. E1 confirmed that this applicant should not be approved and corrected its label to “rejected”.
Subsequently, E1 examined the rule with the second-highest anomaly score. This rule suggested approval for applicants with primarily negative or neutral conditions such as some debt, low credit score, having a driver’s license (Fig. 9A2). E1 doubted the approval decision for these applicants, and checked the samples covered by this rule. He found most of the applicants had no prior default records and a certain level of income (Fig. 9B2). E1 believed these positive conditions were the real reason for the approval. By filtering samples with such positive conditions using the scented widgets in the attribute view (Fig. 9C2), he found that most of them (98.26%) are approved (Fig. 9D2). Thus, he suspected the occurrence of this rule is because of the spurious association between the irrelevant conditions and the approval decision in these covered samples (Fig. 9A2). Discussion with E2 suggested adding more diverse samples, especially those that meet the conditions of this rule but get rejected, to mitigate this spurious association. E2 then generated more samples using data augmentation, and E1 validated these samples to ensure they accurately reflected real-world scenarios.